# MathScape: Evaluating MLLMs in multimodal Math Scenarios through a Hierarchical Benchmark

### v2

@@ 3.1 Construction of MathScape
As illustrated in Figure 3(b), the images are then transformed to closely align with real-world scenarios by capturing photos of printed images, screen displays, and using screenshots.
@@ 4.3 Stability Results and Analysis
Figure 7: The variation of accuracy with answer length.
Figure 8: The variation of accuracy with answer length.
@@ 5.1 Challenges
This issue involves the accurate understanding of the spatial layout of graphics. For instance, in geometry problems, errors can occur if the model fails to correctly recognize the lengths or angles of figures. Such mistakes often stem from a lack of deep understanding of graphic properties or insufficient ability to shift perspectives. As illustrated in Figure 10, Case Study 2, the model incorrectly interprets the distance from point A to 0 as $2\sqrt${2}square-root start_ARG 2 end_ARG.

### v3

@@ 3.1 Construction of MathScape
As illustrated in Figure 3(b), the images are then transformed to closely align with real-world scenarios by capturing photos of printed images and screen displays.
@@ 4.2 Performance of Various Models
Figure 7: The variation of accuracy with answer length.
@@ 4.4 Answer Length and Accuracy
Figure 8: The variation of accuracy with answer length.
@@ 5.1 Challenges
This issue involves the accurate understanding of the spatial layout of graphics. For instance, in geometry problems, errors can occur if the model fails to correctly recognize the lengths or angles of figures. Such mistakes often stem from a lack of deep understanding of graphic properties or insufficient ability to shift perspectives. In Figure 10, Case Study 2, the model incorrectly interprets the distance from point A to 0 as $2\sqrt${2}square-root start_ARG 2 end_ARG.
@@ Appendix C Visualization of MathScape
Appendix D Reproducibility Checklist
Question1: Includes a conceptual outline and/or pseudocode description of AI methods introduced?
Answer: Yes.
Question2: Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results?
Answer: Yes.
Question3: Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper?
Answer: Yes.
Question4: Does this paper make theoretical contributions?
Answer: Yes.
Question5: All assumptions and restrictions are stated clearly and formally?
Answer: Yes.
Question6: All novel claims are stated formally (e.g., in theorem statements)?
Answer: Yes.
Question7: Proofs of all novel claims are included?
Answer: Yes.
Question8: Proof sketches or intuitions are given for complex and/or novel results?
Answer: Yes.
Question9: Appropriate citations to theoretical tools used are given?
Answer: Yes.
Question10: All theoretical claims are demonstrated empirically to hold?
Answer: Yes.
Question11: All experimental code used to eliminate or disprove claims is included?
Answer: Yes.
Question12: Does this paper rely on one or more datasets?
Answer: Yes.
Question13: A motivation is given for why the experiments are conducted on the selected datasets?
Answer: Yes.
Question14: All novel datasets introduced in this paper are included in a data appendix?
Answer: Yes.
Question15: All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question16: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations?
Answer: Yes.
Question17: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available?
Answer: Yes.
Question18: All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing?
Answer: Yes.
Question19: Does this paper include computational experiments?
Answer: Yes.
Question20: Any code required for pre-processing data is included in the appendix?
Answer: Yes.
Question21: All source code required for conducting and analyzing the experiments is included in a code appendix?
Answer: Yes.
Question22: All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question23: All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from?
Answer: Yes.
Question24: If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results?
Answer: Yes.
Question25: This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks?
Answer: Yes.
Question26: This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics?
Answer: Yes.
Question27: This paper states the number of algorithm runs used to compute each reported result?
Answer: Yes.
Question28: Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information?
Answer: Yes.
Question29: The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank)?
Answer: Yes.
Question30: This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments?
Answer: Yes.
Question31: This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting?
Answer: Yes.